# Eigenstructure of the linearized electrical impedance tomography problem under radial perturbations

### v1

@@ 1. Introduction
While standard theory guarantees the continuity of $F:L^{\infty}(B)\to\mathscr{L}(L^{2}_{\diamond}({\partial B}))$, it offers little insight into its approximation by finite-rank operators. According to [5, Appendix A], the mapping $F$ continuously extends to perturbations on $L^{d}(B)$, but its approximability by finite-rank operators and compactness remain unstudied. In this work, we show that the Fréchet derivative $F$ considered in a stronger topology, more specifically as a mapping $F:L^{2}(B)\to\mathscr{L}(L^{2}_{\diamond}({\partial B}))$ is approximable by finite-rank operators, whenever the perturbations $\eta\in L^{2}(B)$ are rotationally symmetric.
To the best of our knowledge, the extension to $L^{2}(B)$ perturbations has only been shown in two spatial dimension (see the article by Garde and Hyvönen [6]), here we establish the extension in any spatial dimension $d\geq 2$ for any $\eta$ in an infinite dimensional subspace of $L^{2}(B)$. We believe that this is the first result proving that the Fréchet derivative is compact and approximable by finite-rank operators with respect to a topology stronger (or equally strong) than the standard $\mathscr{L}(L^{\infty}(B),\mathscr{L}(L^{2}_{\diamond}({\partial B})))$ topology. The extension to $L^{2}(B)$ perturbations and the approximability by finite-rank operators are desired properties from the standpoint of numerical implementations.
This paper is organized as follows. Section 2 presents the main results, including Theorem 2.1, which characterizes the eigenstructure of the linear approximation $F\eta$ for $\eta$ belonging to an infinite-dimensional subspace of $L^{2}(B)$, and Corollary 2.2, which establishes the compactness of $F$ when restricted to the same subspace. Section 3 introduces two auxiliary results, Lemma 3.1 and Lemma 3.2, which serve as key tools in the proof of the main theorem. The proof of Theorem 2.1 is then given in Section 4.
@@ 2. The main results
A key ingredient in this argument is a general result from Banach space theory: given Banach spaces $X$ and $Y$, a linear and bounded operator $A\in\mathscr{L}(X,Y)$ is compact if it can be approximated in the operator norm by a sequence of finite-rank operators $A_{n}$. This follows from the fact that the space of compact linear operators $\mathscr{K}(X,Y)\subset\mathscr{L}(X,Y)$ is closed in the operator topology, and that every finite-rank operator is compact.
@@
In an earlier study [7, Section 6.1], the radial singular functions of truncated $F$ were computed numerically. At the time, these functions were not explicitly interpreted as approximations for singular functions of $F$ itself. However, Corollary 2.2 suggests that repeating the numerical experiment within the framework and assumptions of the present work would yield approximations of the radial singular vectors of $F$.
@@ 3.1. Spherical harmonics
We start by recalling some facts about spherical harmonics; for additional insights on the topic, consult for example [8, Chapter 5] and [9].
@@ 3.2. Jacobi polynomials
The definition of (3.5) is in accordance with [10, §18.5(iii) Eq. 18.5.7], after using a change of variables $x\mapsto 1-2r$, fixing the parameters $\alpha=d-1$ and $\beta=0$, and normalizing with respect to the $L^{2}_{r^{d-1}((0,1))}$ inner product. With these modification, we have orthonormality with respect to the order index $k$,
@@
We denote the Kronecker delta by $\delta_{k,k^{\prime}}$ in equation (3.6). Correspondingly, the recurrence identity [10, §18.9(i) Eq. 18.9.2_1] takes the form
@@
We begin the proof by assuming that $\eta\in\mathcal{R}$ is fixed and define the following series expansion
@@
We observe that, in order for the inner product to be potentially non-zero, the indices of the spherical harmonics must coincide.
@@
$\leq\lvert{\partial B}\rvert^{-\frac{1}{2}}\|\eta\|_{L^{2}(B)}\,\sqrt{\sum^{2\ell-2}_{k=0}\frac{2k+d}{\ell^{2}}\frac{(2\ell-2+d)!\,(2\ell-2)!}{(2\ell-2+d+k)!\,(2\ell-2-k)!}}.$
@@
Next, we show that the sum
$\sum^{2\ell-2}_{k=0}\frac{2k+d}{\ell^{2}}\frac{(2\ell-2+d)!\,(2\ell-2)!}{(2\ell-2+d+k)!\,(2\ell-2-k)!}$ (4.7)
behaves as $\mathcal{O}(\ell^{-1})$ when $\ell\to\infty$. Since we are unable to derive a closed-form expression for (4.7), and its convergence to zero is not immediately evident in its current form, we resort to an integral test to study the behavior as a function of $\ell$. Before proceeding to the integral test, let us construct an upper bound for the fraction of factorials in (4.7), which is undoubtedly the most tedious part of the summand. To achieve this, we employ a technique developed in an earlier work [11].
Let us denote $L=2\ell-1$ and define a dummy function
$\rho(x):=\frac{\Gamma(L+d)\,\Gamma(L)}{\Gamma(L+d+x)\,\Gamma(L-x)},\quad 0\leq x\leq L-1,$ (4.8)
which is the continuous version of the fraction of factorials in (4.7) with the index of summation $k$ replaced by a continuous variable $x$. Also, we adopt the usual definition of the gamma function, which means that $\Gamma(n)=(n-1)!$.
We may differentiate (4.8) with respect to the variable $x$. After using the product rule and the identity
$\frac{\mathrm{d}}{\mathrm{d}z}\frac{1}{\Gamma(z)}=-\frac{\psi(z)}{\Gamma(z)},$
where $\psi(z)$ is the digamma function, we get
$\frac{\mathrm{d}}{\mathrm{d}x}\rho(x)=\rho(x)\big(\psi(L-x)-\psi(L+d+x)\big).$ (4.9)
The digamma function is strictly monotonically increasing on the positive real axis, and thus the derivative (4.9) is negative for all $0\leq x\leq L-1$.
Our strategy is as follows: we replace the difference of digamma functions in (4.9) with a less negative expression, and then resort to Grönwall’s inequality in order to construct an upper bound for $\rho$. The hope is that the constructed upper bound simplifies the analysis of (4.7).
The derivative of the digamma function is the trigamma function $\psi_{1}(z)$ which is a strictly convex and strictly decreasing function on the positive real axis. According to [12, Lemma 1], we have a lower bound
$\psi_{1}(z)\geq z^{-1},\qquad z>0.$ (4.10)
We may employ the fundamental theorem of calculus, Jensen’s inequality and (4.10) to get an estimate for the difference of digamma functions in (4.9):
$-\left(\psi(L+d+x)-\psi(L-x)\right)$ $=-(2x+d)\bigg(\frac{1}{2x+d}\int^{L+d+x}_{L-x}\psi_{1}(z)\,\mathrm{d}z\bigg)$
$\leq-(2x+d)\,\psi_{1}\bigg(\frac{1}{2x+d}\int^{L+d+x}_{L-x}z\,\mathrm{d}z\bigg)$
$\leq-\frac{2x+d}{L+\frac{d}{2}}.$
Since $\rho(x)>0$ for all $0\leq x\leq L-1$, we have
$\frac{\mathrm{d}}{\mathrm{d}x}\rho(x)\leq\rho(x)\bigg(-\frac{2x+d}{L+\frac{d}{2}}\bigg),$
and from the differential form of Grönwall’s inequality, a simple upper bound is acquired,
$\rho(x)\leq\rho(0)\exp\bigg(-\int_{0}^{x}\frac{2z+d}{L+\frac{d}{2}}\mathrm{d}z\bigg)={\rm e}^{-\frac{2x(x+d)}{2L+d}},\qquad 0\leq x\leq L-1.$ (4.11)
Collecting the various estimates together and returning to the discrete setting, we obtain
$\lvert\lambda_{\ell}(\eta)\rvert\leq\lvert{\partial B}\rvert^{-\frac{1}{2}}\|\eta\|_{L^{2}(B)}\,\sqrt{\sum^{2\ell-2}_{k=0}\frac{2k+d}{\ell^{2}}\,\mathrm{e}^{-\frac{2k(k+d)}{4\ell-2+d}}}.$ (4.12)
The summand in (4.12) is increasing as a function of $k$ on the interval $(0,k^{*})$ and decreasing on the interval $(k^{*},2\ell-2)$ with respect to a unique critical point
$k^{*}=\tfrac{1}{2}\big(\sqrt{4\ell-2+d}-d\big).$
Therefore, in the integral test the maximal value needs to be taken into account, and we may estimate
$\sum^{2\ell-2}_{k=0}\frac{2k+d}{\ell^{2}}\frac{(2\ell-2+d)!\,(2\ell-2)!}{(2\ell-2+d+k)!\,(2\ell-2-k)!}$ $\leq\sum^{2\ell-2}_{k=0}\frac{2k+d}{\ell^{2}}\,\mathrm{e}^{-\frac{2k(k+d)}{4\ell-2+d}}$
$\leq\int^{2\ell-2}_{0}\frac{2x+d}{\ell^{2}}\,\mathrm{e}^{-\frac{2x(x+d)}{4\ell-2+d}}\,\mathrm{d}x+\frac{2k^{*}+d}{\ell^{2}}\,\mathrm{e}^{-\frac{2k^{*}(k^{*}+d)}{4\ell-2+d}}$
$\leq\int^{2\ell-2}_{0}\frac{2x+d}{\ell^{2}}\,\mathrm{e}^{-\frac{2x(x+d)}{4\ell-2+d}}\,\mathrm{d}x+\frac{\sqrt{4\ell-2+d}}{\ell^{2}}\,\mathrm{e}^{d}.$
For the integral, we first perform a change of variables $x\mapsto\sqrt{\ell}\,z$, then expand the domain of integration, and finally eliminate the $\ell$ dependence from the fraction and the exponent:
$\int^{2\ell-2}_{0}\frac{2x+d}{\ell^{2}}\,\mathrm{e}^{-\frac{2x(x+d)}{4\ell-2+d}}\,\mathrm{d}x$ $=\int^{2\sqrt{\ell}-\frac{2}{\sqrt{\ell}}}_{0}\frac{2\sqrt{\ell}z+d}{\ell^{\frac{3}{2}}}\,\mathrm{e}^{-\frac{2\sqrt{\ell}z(\sqrt{\ell}z+d)}{4\ell-2+d}}\,\mathrm{d}z$
$\leq\int^{\infty}_{0}\frac{2\sqrt{\ell}z+d}{\ell^{\frac{3}{2}}}\,\mathrm{e}^{-\frac{2\ell z^{2}}{4\ell-2+d}}\,\mathrm{d}z$
$\leq\ell^{-1}\int^{\infty}_{0}(2z+d)\,\mathrm{e}^{-\frac{2z^{2}}{d+4}}\,\mathrm{d}z=\ell^{-1}\,\left(\frac{d(\sqrt{2\pi}\sqrt{d+4}+2)}{4}+2\right).$
Combining the estimate for the maximal value and for the integral leads to the following conservative upper bound
$\sum^{2\ell-2}_{k=0}\frac{2k+d}{\ell^{2}}\frac{(2\ell-2+d)!\,(2\ell-2)!}{(2\ell-2+d+k)!\,(2\ell-2-k)!}$ $\leq\ell^{-1}\left(\frac{d(\sqrt{2\pi}\sqrt{d+4}+2)}{4}+2\right)+\frac{\sqrt{4\ell-2+d}}{\ell^{2}}\,\mathrm{e}^{d}$
$\leq 4\,d^{2}\,\mathrm{e}^{d}\ell^{-1}.$
@@
Finally, we may substitute this estimate into (4.12) and apply the well-known formula
@@
$\lvert\lambda_{\ell}(\eta)\rvert\leq C_{d}\,\|\eta\|_{L^{2}(B)}\,\ell^{-\frac{1}{2}},\quad C_{d}=d\,\big(\tfrac{\mathrm{e}^{2}}{\pi}\big)^{\frac{d}{4}}\,\sqrt{2\,\Gamma\big(\tfrac{d}{2}\big)}.$ (4.13)
Although the constant $C_{d}$ may be suboptimal, we prefer it for its simplicity.
@@ Acknowledgements
I wish to thank Nuutti Hyvönen for insightful discussions and guidance related to the subject of the paper. The project is supported by the Research Council of Finland (decisions 353081 and 359181).

### v2

@@ 1. Introduction
While standard theory guarantees the continuity of $F:L^{\infty}(B)\to\mathscr{L}(L^{2}_{\diamond}({\partial B}))$, it offers little insight into its approximation by finite-rank operators. According to [5, Appendix A], the mapping $F$ continuously extends to perturbations on $L^{d}(B)$, but its approximability by finite-rank operators and compactness remain unstudied. In this work, we consider the Fréchet derivative $F$ on the subspace $\mathcal{R}\subset L^{2}(B)$ consisting of rotationally symmetric perturbations (cf. (2.2)). We then study $F$ in a stronger topology, more specifically as a mapping $F:\mathcal{R}\to\mathscr{L}(L^{2}_{\diamond}({\partial B}))$, and prove that $F$ can be approximated by finite-rank operators.
To the best of our knowledge, the extension to $L^{2}(B)$ perturbations has only been shown in two spatial dimension (see the article by Garde and Hyvönen [6]), here we establish the extension in any spatial dimension $d\geq 2$ for any $\eta$ in an infinite dimensional subspace $\mathcal{R}\subset L^{2}(B)$. We believe that this is the first result proving that the Fréchet derivative is compact and approximable by finite-rank operators with respect to a topology stronger (or equally strong) than the standard $\mathscr{L}(L^{\infty}(B),\mathscr{L}(L^{2}_{\diamond}({\partial B})))$ topology. The extension to $\mathcal{R}$ perturbations and the approximability by finite-rank operators are desired properties from the standpoint of numerical implementations.
The (linearized) Calderón problem in the radial setting has recently been studied in the context of Born approximations [7, 8]. Closely related to this approach, the reader may also find the recent preprints [9, 10] of interest.
1.1. Article structure
In summary, the remainder of this text is structured as follows. Section 2 presents the main results, including Theorem 2.1, which characterizes the eigenstructure of the linear approximation $F\eta$ for $\eta$ belonging to an infinite-dimensional subspace of $L^{2}(B)$, and Corollary 2.2, which establishes the compactness of $F$ when restricted to the same subspace. Section 3 introduces two auxiliary results, Lemma 3.1 and Lemma 3.2, which serve as key tools in the proof of the main theorem. Section 4 contains the proof of Theorem 2.1 and the auxiliary result Lemma 4.1.
@@ 2. The main results
A key ingredient in this argument is a general result from Banach space theory: given Banach spaces $X$ and $Y$, a linear and bounded operator $A\in\mathscr{L}(X,Y)$ is compact if it can be approximated in the operator norm by a sequence of finite-rank operators. Here, the operator norm is defined in the usual way by
$\|A\|_{\rm op}:=\sup_{x\in X,\,\|x\|_{X}=1}\|Ax\|_{Y}.$
This result follows from the fact that the space of compact operators $\mathscr{K}(X,Y)\subset\mathscr{L}(X,Y)$ is closed in the operator topology, and that every finite-rank operator is compact.
@@
In an earlier study [11, Section 6.1], the radial singular functions of truncated $F$ were computed numerically. At the time, these functions were not explicitly interpreted as approximations for singular functions of $F$ itself. However, Corollary 2.2 suggests that repeating the numerical experiment within the framework and assumptions of the present work would yield approximations of the radial singular vectors of $F$.
@@ 3.1. Spherical harmonics
We start by recalling some facts about spherical harmonics; for additional insights on the topic, consult for example [12, Chapter 5] and [13].
@@ 3.2. Jacobi polynomials
The definition of (3.5) is in accordance with [14, §18.5(iii) Eq. 18.5.7], after using a change of variables $x\mapsto 1-2r$, fixing the parameters $\alpha=d-1$ and $\beta=0$, and normalizing with respect to the $L^{2}_{r^{d-1}((0,1))}$ inner product. With these modification, we have orthonormality with respect to the order index $k$,
@@
We denote the Kronecker delta by $\delta_{k,k^{\prime}}$ in equation (3.6). Correspondingly, the recurrence identity [14, §18.9(i) Eq. 18.9.2_1] takes the form
@@
We begin the proof by assuming that $\eta\in\mathcal{R}$ is fixed and perform the following series expansion
@@
We observe that, in order for the inner product to be potentially non-zero, the indices of the spherical harmonics must coincide. Note that Lemma 3.1 is applicable in (4.3) only because $\eta$ has essentially no angular component. Indeed, even a mild angular dependence of $\eta$ would prevent Lemma 3.1 from being applied in its current form.
@@
$\leq\lvert{\partial B}\rvert^{-\frac{1}{2}}\|\eta\|_{L^{2}(B)}\,\sqrt{\sum^{2\ell-2}_{k=0}\frac{2k+d}{\ell^{2}}\frac{(2\ell-2+d)!\,(2\ell-2)!}{(2\ell-2+d+k)!\,(2\ell-2-k)!}}.$ (4.7)
@@
Before proceeding with the proof, let us derive an auxiliary result in the form of a lemma.
Lemma 4.1.
Let $L\in\mathbb{N}_{0}$ and $d\in\mathbb{R}$, then
$\sum^{L}_{k=0}(2k+d)\frac{\Gamma(s+1)}{\Gamma(k+s+1)}\frac{L!}{(L-k)!}=L+d.$ (4.8)
Proof.
Let $L\in\mathbb{N}_{0}$ and $d\in\mathbb{R}$, and denote $s:=L+d$. The case $L=0$ is straightforward, so in what follows we restrict our attention to $L\geq 1$, where the argument is nontrivial.
For $k=0$ the summand in (4.8) reduces to $d$, so we can write
$\mu=d+\sum^{L}_{k=1}(2k+d)\frac{\Gamma(s+1)}{\Gamma(k+s+1)}\frac{L!}{(L-k)!}.$ (4.9)
The goal is to show that $\mu=d+L$.
We start by recalling the binomial theorem, which for $(1+y)^{L}$ reads
$\sum^{L}_{k=0}\binom{L}{k}y^{k}=(1+y)^{L},\quad y\in\mathbb{R}.$ (4.10)
We will use (4.10) to derive two identities used later in this proof. The first identity is acquired by differentiating (4.10) with respect to $y$, which results in
$\sum^{L}_{k=1}k\binom{L}{k}y^{k-1}=L(1+y)^{L-1}.$ (4.11)
The second identity is obtained by differentiating (4.10) twice, multiplying both sides with $y$ and after that adding (4.11) on both sides:
$\left(\sum^{L}_{k=2}k(k-1)\binom{L}{k}y^{k-2}\right)y+\sum^{L}_{k=1}k\binom{L}{k}y^{k-1}=\left(L(L-1)(1+y)^{L-2}\right)y+L(1+y)^{L-1},$
which simplifies to
$\sum^{L}_{k=1}k^{2}\binom{L}{k}y^{k-1}=L(1+Ly)(1+y)^{L-2}.$ (4.12)
The next step is to use the beta function and write the to-be evaluated sum as an integral. The integral form of the beta function for $z_{1},z_{2}\in\mathbb{C}$, with positive real parts, i.e., $\Re(z_{1}),\Re(z_{2})>0$ is
$B(z_{1},z_{2})=\frac{\Gamma(z_{1})\Gamma(z_{2})}{\Gamma(z_{1}+z_{2})}=\int^{1}_{0}t^{z_{1}-1}(1-t)^{z_{2}-1}\,\mathrm{d}t.$
Thus, for the summation index $k\geq 1$ we may write
$\frac{\Gamma(s+1)}{\Gamma(k+s+1)}=\frac{1}{(k-1)!}\int^{1}_{0}t^{s}(1-t)^{k-1}\,\mathrm{d}t.$ (4.13)
Defining
$\xi(t):=\sum^{L}_{k=1}(2k+d)k\binom{L}{k}(1-t)^{k-1},$ (4.14)
it follows from (4.13) that
$\mu=d+\int^{1}_{0}t^{s}\xi(t)\,\mathrm{d}t.$ (4.15)
We now apply the identities (4.11) and (4.12) to derive a closed-form expression for $\xi(t)$. To this end, we substitute y=(1t)y=(1-t) in (4.11) and (4.12), and form a linear combination of the resulting expressions. This yields
$\xi(t)$ $=2\left(L\big(1+L(1-t)\big)\big(1+(1-t)\big)^{L-2}\right)+d\left(L\big(1+(1-t)\big)^{L-1}\right)$
$=L(2-t)^{L-2}(2+2s-Lt-st).$ (4.16)
It is straightforward to verify that
$t^{s}\xi(t)=\frac{\mathrm{d}}{\mathrm{d}t}\left(Lt^{s+1}(2-t)^{L-1}\right).$
Evaluating the integral (4.15) using the above identity yields $\mu=d+L$, completing the proof. ∎
We now proceed with the proof of Theorem 2.1 and apply Lemma 4.1 to evaluate the sum in (4) exactly:
$\sum^{2\ell-2}_{k=0}\frac{2k+d}{\ell^{2}}\frac{(2\ell-2+d)!\,(2\ell-2)!}{(2\ell-2+d+k)!\,(2\ell-2-k)!}=\frac{2\ell-2+d}{\ell^{2}}.$ (4.17)
Since, $\ell\geq 1$ we further estimate:
$\frac{2\ell-2+d}{\ell^{2}}\leq\frac{d+2}{\ell}.$ (4.18)
@@
Finally, we may substitute this estimate into (4) and apply the well-known formula
@@
$\lvert\lambda_{\ell}(\eta)\rvert\leq C_{d}\,\|\eta\|_{L^{2}(B)}\,\ell^{-\frac{1}{2}},\quad C_{d}=\sqrt{\frac{(d+2)\,\Gamma\big(\frac{d}{2}\big)}{2\pi^{\frac{d}{2}}}}.$ (4.19)
@@ Acknowledgements
I wish to thank Nuutti Hyvönen for insightful discussions and guidance related to the subject of the paper, and I wish to thank Vanni Noferini for providing a proof for Lemma 4.1. The project is supported by the Research Council of Finland (decisions 353081 and 359181).